OpenStudy (anonymous):

One number is 11 more than twice another number. If the sum of the numbers is twice their difference, find the numbers.

OpenStudy (anonymous):

a=2b+11 a+b = 2(a-b) now solve a+b = 2a-2b 3b = a, put that into the 1st equation: 3b = 2b+11 b=11 a = 2(11) +11 = 33 check in the 2nd equation: 33+11 = 2(33-11) 44 = 2(22) 44 = 44

OpenStudy (anonymous):

thanks